# Gamma-ray spectra of the Crab, Vela and Geminga pulsars fitted with SED of the emission from their current sheet

### 24 Jan 2024 v2

@@ 3.2 Spectrum of the Vela pulsar (PSR J0835-4510)
The data on the Vela pulsar both in the GeV band detected by Fermi-LAT (Abdo et al., 2013) and in the TeV band detected by H.E.S.S. (The H.E.S.S. Collaboration et al., 2023) are shown in Fig. 2. The SED that best fits these data (represented by the red curve in Fig. 2) is described by the expression in Eq. (5) for the following values of its free parameters:
@@
$\sigma_{0}$ $=$ $(3.12\pm 1.70)\times 10^{-8},\quad\zeta=0.518\pm 0.065,$
@@
The difference between the values of the free parameters in the above two ranges of photon energies stems from the fact that the higher-frequency radiation is appreciably more focused than its lower-frequency counterpart in this case. This is reflected not only in the much lower value of $\sigma_{0}$ in Eq. (21) compared to that in Eq. (20), but also in the higher value of $\rho$ (i.e. the shorter range of values of $\sigma_{21}$): the smaller the value of $\sigma_{21}$, the closer to each other are the stationary points of the phases of the emitted waves and so the more focused is the radiation (Ardavan, 2021, Sect. 4.5).
Figure 3: Gamma-ray spectrum of the Geminga pulsar. The FERMI-LAT data points below $35$ GeV (coloured black) are those reported in Abdo et al. (2013). The MAGIC data points above $35$ GeV (coloured blue) are those reported in The MAGIC Collaboration et al. (2020). The curve (coloured red) is a plot of the SED described by Eq. (5) for the parameters given in Eqs. (22) and (23).
@@ 3.3 Spectrum of the Geminga pulsar (PSR J0633+1746)
The Fermi-LAT (Abdo et al., 2013) and MAGIC (The MAGIC Collaboration et al., 2020) data on the gamma-ray emission from the Geminga pulsar are shown in Fig. 3. The fit to these data (the red curve in Fig 3) is described by the expression in Eq. (5) for
@@
$\sigma_{0}$ $=$ $(2.17\pm 0.99)\times 10^{-8},\quad\zeta=0.646\pm 0.028$
@@
From Eqs. (6), (4) and (1), it follows that the parameters $\kappa$, $\zeta$, $\rho$ and $\sigma_{0}$ in Eq. (5) are related to the characteristics of the source of the observed radiation via the quantities $\omega$, $\kappa_{0}$, $\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}^{(2)}$ that appear in the expression for the SED ${\cal S}_{\nu}$. In this section we use the results of the analysis presented in Ardavan (2021) to express these quantities in terms of the inclination angle of the central neutron star, $\alpha$, the magnitude of the star’s magnetic field at its magnetic pole, $B_{0}=10^{12}{\hat{B}}_{0}$ Gauss, the radius of the star, $r_{s0}=10^{6}d$ cm, the rotation frequency of the star, $\omega=10^{2}{\hat{P}}^{-1}$ rad/s, and the spherical polar coordinates, $R_{P}=D\,\,{\rm kpc}=3.085\times 10^{21}D$ cm, $\varphi_{P}$ and $\theta_{P}$, of the observation point $P$ in a frame whose centre and $z$-axis coincide with the centre and spin axis of the star.
When $\sigma_{21}\ll 1$ and so the quantities $\tau_{2{\rm min}}$ and $\tau_{2{\rm max}}$ that appear in Eqs. (177), (138)–(146) and (136) of Ardavan (2021) are approximately equal, these equations imply that
@@
$a$ $=$ $({\hat{r}}_{P}^{2}{\hat{r}}_{sC}^{2}\sin^{2}\theta-1)\left[\frac{({\hat{r}}_{P}^{2}{\hat{r}}_{sC}^{2}\sin^{2}\theta-1)^{1/2}+{\hat{r}}_{sC}}{{\hat{r}}_{sC}({\hat{r}}_{P}^{2}-1)^{1/2}({\hat{R}}_{P}^{2}\sin^{2}\theta-1)^{1/2}}\right.$ (28)
$\left.-\frac{1}{({\hat{r}}_{P}^{2}{\hat{r}}_{sC}^{2}\sin^{2}\theta-1)^{1/2}}\right],$
@@
$\theta=\arccos(\sin\alpha\cos\tau),$ (32)
$\left|{\tilde{\bf P}}_{2}\right|$ $=$ $\cos\alpha[{\hat{r}}_{sC}^{2}(1+\cos^{2}\theta_{P}-2\cos\theta\cos\theta_{P})+\tan^{2}\alpha$
$-\cot^{2}\theta+2{\hat{r}}_{sC}\sin\theta\cos\theta_{P}(\tan^{2}\alpha-\cot^{2}\theta)^{1/2}]^{1/2},\quad\,\,$
@@
and the variables $\tau_{2{\rm min}}$ and $\tau_{2{\rm max}}$ stand for the minimum and maximum of the function $f_{2C}$ (see also Eqs. (7), (9), (174), (175), (93)–(95), (97) and (88) of Ardavan 2021). The caret on $R_{P}$ and $r_{s0}$ (and $r_{P}=R_{P}\sin\theta_{P}$, $z_{P}=R_{P}\cos\theta_{P}$) is used here to designate a variable that is rendered dimensionless by being measured in units of the light-cylinder radius $c/\omega$. (Note the following two corrections: the vector ${\bf P}_{2}$ in Eq. (145) and the numerical coefficient $1.54\times 10^{19}$ in Eq. (177) of Ardavan 2021 have been corrected to read ${\tilde{\bf P}}_{2}$ and $4.15\times 10^{18}$, respectively.)
@@
The expression for $|{\tilde{\bf P}}_{2}|$ in Eq. (LABEL:E24) is derived from that for ${\bf P}_{2}$ given by Eqs. (98), (80), (78) and (62) of Ardavan (2021). In this derivation, we have set the observation point on the cusp locus of the bifurcation surface where ${\hat{r}}_{s}={\hat{r}}_{sC}$, have approximated $(p_{1},p_{2},p_{3})$ by its far-field value $2^{1/3}{\hat{R}}_{P}^{-1}({\hat{R}}_{P}^{-1},-1,1)$ and have let ${\bf P}_{2}={\hat{R}}_{P}^{-1}{\tilde{\bf P}}_{2}$. The factor ${\hat{R}}_{P}^{-2}$ that would have otherwise appeared in the resulting expression for $|\boldsymbol{\cal P}^{(2)}_{2}|^{2}$ is thus incorporated in the factor $\kappa_{0}$ in Eq. (24).
@@
For certain values of $\theta_{P}$, denoted by $\theta_{P2S}$, the function $f_{2C}(\tau)$ has an inflection point (see Ardavan 2021, Sect. 4.4, and Ardavan 2023, Sect. 2). For any given inclination angle $\alpha$, the position $\tau_{2S}$ of this inflection point and the colatitude $\theta_{P2S}$ of the observation points for which $f_{2C}(\tau)$ has an inflection point follow from the solutions to the simultaneous equations $\partial f_{2C}/\partial\tau=0$ and $\partial^{2}f_{2C}/\partial\tau^{2}=0$. (Explicit expressions for the derivatives that appear in these equations can be found in Appendix A of Ardavan 2021.) For values of $\theta_{P}$ close to $\theta_{P2S}$, the separation between the maximum $\tau=\tau_{2{\rm max}}$ and minimum $\tau=\tau_{2{\rm min}}$ of $f_{2C}$ and hence the value of $\sigma_{21}$ are vanishingly small. Since the high-frequency radiation that arises from the current sheet is detectable only in the vicinity of the latitudinal direction $\theta_{P}=\theta_{P2S}$ (see Fig. 4), here we have evaluated the variables that appear in Eqs. (25)–(LABEL:E24) for $\sigma_{21}\ll 1$. The ratio appearing inside the second pair of parentheses in Eq. (25), which becomes indeterminate as $\sigma_{21}$ tends to zero, approaches a finite value in this limit.
@@
The dimensionless frequency $k=2\pi\nu/\omega$ of a $1$ GeV photon has the value $8\times 10^{21}$ in the case of the Crab pulsar. In contrast, as indicated by the small values of the fit parameter $\sigma_{0}$ encountered in Sect. 3, the value of the variable $\chi=2\sigma_{0}^{3}k/3$ that appears in the description of the SED ${\cal S}_{\nu}$ is of the order of $10^{-3}$ for a $1$ GeV gamma-ray photon. Because any measurement of the spectrum described by Eq. (5) is carried out by counting the number of detected gamma-ray photons in a bin covering an energy interval of a few MeV or GeV, the widths of the equivalent frequency bins involved in such measurements are of the order of $\Delta k=3\sigma_{0}^{-3}\Delta\chi/2$ with $10^{-4}\lesssim\Delta\chi\lesssim 10^{-2}$.
@@
While ${\hat{\kappa}}_{\rm obs}$ only contains the observed parameters of the pulsar and its emission, the value of ${\hat{\kappa}}_{\rm th}$ is determined by the physical characteristics of the magnetospheric current sheet that acts as the source of the observed emission. When, as in Fig. 20 of Ardavan (2021), the value of the colatitude $\theta_{P}$ of the observation point is sufficiently close to that of the critical angle $\theta_{P2S}$ for $|\tau_{2{\rm max}}-\tau_{2{\rm min}}|$ and hence $\sigma_{21}$ to be much smaller than unity, the right-hand side of Eq. (35) is a function of the inclination angle $\alpha$ and the observer’s distance ${\hat{R}}_{P}$ only.
Equation (34) thus enables us to connect the the parameters of the fitted spectra to the physical characteristics of their sources.
@@ 4.1 Characteristics of the source of the Crab pulsar’s emission
for $\Delta\chi=10^{-3}$. The dependence of ${\hat{B}}_{0}d^{2}$ on $\alpha$ implied by this relationship, in conjunction with Eq. (35), is plotted in Fig. 5.
One can infer from the separation between the two peaks in the gamma-ray light curve of the Crab pulsar  ($\sim 0.4$ cycles; Abdo et al., 2010) that the inclination angle $\alpha$ of this pulsar has a value lying in the range $60^{\circ}$ to $70^{\circ}$ (Ardavan, 2021, Sect. 5.4). This in turn implies that the value of ${\hat{B}}_{0}d^{2}$ lies between $1.05$ and $2.61$ in this case (see Fig. 5), i.e. that the magnitude of the star’s magnetic field at its magnetic pole lies in the range $1.05\times 10^{12}$ to $2.61\times 10^{12}$ Gauss if we assume that the star’s radius has the value $10^{6}$ cm. These are of the same order of magnitude as the value of $B_{0}$ obtained from the conventional formula for magnetic dipole radiation (i.e. $3.79\times 10^{12}$ Gauss; Manchester et al. 2005).
The corresponding value of the privileged latitudinal direction $\theta_{P2S}$ in which the high-frequency radiation from the Crab pulsar can be observed lies between $90^{\circ}$ and $117^{\circ}$ according to Fig. 4. The range of values of the inclination angle $\alpha$ determines also the scale factor $j_{0}$ of the electric current density in the magnetosphere of the pulsar: from Eq. (22) of Ardavan (2021) it follows that
@@
The above results supersede those reported in Ardavan (2023, Sect. 4).
@@ 4.2 Characteristics of the source of the Vela pulsar’s emission
Properties of the central neutron star of the Vela pulsar and its magnetosphere are more accurately reflected in the parameters of its spectrum over the photon energies $0.1$ to $10$ GeV, i.e. the interval that encompasses the majority of the data points plotted in Fig. 2. Insertion of the value of the fit parameter $\kappa$ given in Eq. (20), the period ($0.089$ s) and the distance ($287$ pc) of this pulsar in Eq. (36), together with ${\hat{\kappa}}_{\rm obs}={\hat{\kappa}}_{\rm th}$, results in
@@
for $\Delta\chi=10^{-3}$. Thus the dependence of ${\hat{B}}_{0}d^{2}$ on $\alpha$ is given by a shifted version of the curve in Fig. 5: one that is vertically lowered by 0.109 (cf. Eqs. (37) and (39)).
@@
That the main peaks of the gamma-ray light curve of the Vela pulsar are separated by $\sim 0.43$ cycles (The H.E.S.S. Collaboration et al., 2023) is consistent with an inclination angle close to $70^{\circ}$ (see Ardavan, 2021, Sect. 5.4).

### 03 Mar 2024 v3

@@ 3.2 Spectrum of the Vela pulsar (PSR J0835-4510)
The data on the Vela pulsar both in the GeV band detected by Fermi-LAT [19] and in the TeV band detected by HESS [2] are shown in figure 2. The SED that best fits these data (represented by the red curve in figure 2) is described by the expression in eq. (5) for the following values of its free parameters:
@@
$\sigma_{0}$ $=$ $(3.12\pm 0.18)\times 10^{-8},\quad\zeta=0.518\pm 0.065,$
@@
The difference between the values of the free parameters in the above two ranges of photon energies stems from the fact that the higher-frequency radiation is appreciably more focused than its lower-frequency counterpart in this case. This is reflected not only in the much lower value of $\sigma_{0}$ in eq. (20) compared to that in eq. (19), but also in the higher value of $\rho$ (i.e. the shorter range of values of $\sigma_{21}$): the smaller the value of $\sigma_{21}$, the closer to each other are the stationary points of the phases of the emitted waves and so the more focused is the radiation (see section 4.5 of [3]).
Figure 3: Gamma-ray spectrum of the Geminga pulsar. The Fermi-LAT data points below $35$ GeV (coloured black) are those reported in [19]. The MAGIC data points above $35$ GeV (coloured blue) are those reported in [20]. The curve (coloured red) is a plot of the SED described by eq. (5) for the parameters given in eqs. (21) and (22).
@@ 3.3 Spectrum of the Geminga pulsar (PSR J0633+1746)
The Fermi-LAT [19] and MAGIC [20] data on the gamma-ray emission from the Geminga pulsar are shown in figure 3. The fit to these data (the red curve in figure 3) is described by the expression in eq. (5) for
@@
$\sigma_{0}$ $=$ $(2.17\pm 0.07)\times 10^{-8},\quad\zeta=0.646\pm 0.028$
@@
From eqs. (6), (4) and (1), it follows that the parameters $\kappa$, $\zeta$, $\rho$ and $\sigma_{0}$ in eq. (5) are related to the characteristics of the source of the observed radiation via the quantities $\omega$, $\kappa_{0}$, $\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}^{(2)}$ that appear in the expression for the SED ${\cal S}_{\nu}$. In this section we use the results of the analysis presented in [3] to express these quantities in terms of the inclination angle of the central neutron star, $\alpha$, the magnitude of the star’s magnetic field at its magnetic pole, $B_{0}=10^{12}{\hat{B}}_{0}$ Gauss, the radius of the star, $r_{s0}=10^{6}d$ cm, the rotation frequency of the star, $\omega=10^{2}{\hat{P}}^{-1}$ rad/s, and the spherical polar coordinates, $R_{P}=D\,\,{\rm kpc}=3.085\times 10^{21}D$ cm, $\varphi_{P}$ and $\theta_{P}$, of the observation point $P$ in a frame whose centre and $z$-axis coincide with the centre and spin axis of the star.
When $\sigma_{21}\ll 1$ and so the quantities $\tau_{2{\rm min}}$ and $\tau_{2{\rm max}}$ that appear in eqs. (177), (138)–(146) and (136) of [3] are approximately equal, these equations imply that
@@
$a=({\hat{r}}_{P}^{2}{\hat{r}}_{sC}^{2}\sin^{2}\theta-1)\left[\frac{({\hat{r}}_{P}^{2}{\hat{r}}_{sC}^{2}\sin^{2}\theta-1)^{1/2}+{\hat{r}}_{sC}}{{\hat{r}}_{sC}({\hat{r}}_{P}^{2}-1)^{1/2}({\hat{R}}_{P}^{2}\sin^{2}\theta-1)^{1/2}}-\frac{1}{({\hat{r}}_{P}^{2}{\hat{r}}_{sC}^{2}\sin^{2}\theta-1)^{1/2}}\right],$ (27)
@@
$\theta=\arccos\left(\sin\alpha\cos\tau\right),$ (31)
$\left|{\tilde{\bf P}}_{2}\right|$ $=$ $\cos\alpha[{\hat{r}}_{sC}^{2}(1+\cos^{2}\theta_{P}-2\cos\theta\cos\theta_{P})+\tan^{2}\alpha$ (32)
$-\cot^{2}\theta+2{\hat{r}}_{sC}\sin\theta\cos\theta_{P}(\tan^{2}\alpha-\cot^{2}\theta)^{1/2}]^{1/2},$
@@
and the variables $\tau_{2{\rm min}}$ and $\tau_{2{\rm max}}$ stand for the minimum and maximum of the function $f_{2C}$ (see also eqs. (7), (9), (174), (175), (93)–(95), (97) and (88) of [3]). The caret on $R_{P}$ and $r_{s0}$ (and $r_{P}=R_{P}\sin\theta_{P}$, $z_{P}=R_{P}\cos\theta_{P}$) is used here to designate a variable that is rendered dimensionless by being measured in units of the light-cylinder radius $c/\omega$. (Note the following two corrections: the vector ${\bf P}_{2}$ in eq. (145) and the numerical coefficient $1.54\times 10^{19}$ in eq. (177) of [3] have been corrected to read ${\tilde{\bf P}}_{2}$ and $4.15\times 10^{18}$, respectively.)
@@
The expression for $|{\tilde{\bf P}}_{2}|$ in eq. (32) is derived from that for ${\bf P}_{2}$ given by eqs. (98), (80), (78) and (62) of [3]. In this derivation, we have set the observation point on the cusp locus of the bifurcation surface where ${\hat{r}}_{s}={\hat{r}}_{sC}$, have approximated $(p_{1},p_{2},p_{3})$ by its far-field value $2^{1/3}{\hat{R}}_{P}^{-1}({\hat{R}}_{P}^{-1},-1,1)$ and have let ${\bf P}_{2}={\hat{R}}_{P}^{-1}{\tilde{\bf P}}_{2}$. The factor ${\hat{R}}_{P}^{-2}$ that would have otherwise appeared in the resulting expression for $|\boldsymbol{\cal P}^{(2)}_{2}|^{2}$ is thus incorporated in the factor $\kappa_{0}$ in eq. (23).
@@
For certain values of $\theta_{P}$, denoted by $\theta_{P2S}$, the function $f_{2C}(\tau)$ has an inflection point (see section 4.4 of [3], and section 2 of [12]). For any given inclination angle $\alpha$, the position $\tau_{2S}$ of this inflection point and the colatitude $\theta_{P2S}$ of the observation points for which $f_{2C}(\tau)$ has an inflection point follow from the solutions to the simultaneous equations $\partial f_{2C}/\partial\tau=0$ and $\partial^{2}f_{2C}/\partial\tau^{2}=0$. (Explicit expressions for the derivatives that appear in these equations can be found in appendix A of [3].) For values of $\theta_{P}$ close to $\theta_{P2S}$, the separation between the maximum $\tau=\tau_{2{\rm max}}$ and minimum $\tau=\tau_{2{\rm min}}$ of $f_{2C}$ and hence the value of $\sigma_{21}$ are vanishingly small. Since the high-frequency radiation that arises from the current sheet is detectable only in the vicinity of the latitudinal direction $\theta_{P}=\theta_{P2S}$ (see figure 4), here we have evaluated the variables that appear in eqs. (24)–(32) for $\sigma_{21}\ll 1$. The ratio appearing inside the second pair of parentheses in eq. (24), which becomes indeterminate as $\sigma_{21}$ tends to zero, approaches a finite value in this limit.
@@
The dimensionless frequency $k=2\pi\nu/\omega$ of a $1$ GeV photon has the value $8\times 10^{21}$ in the case of the Crab pulsar. In contrast, as indicated by the small values of the fit parameter $\sigma_{0}$ encountered in section 3, the value of the variable $\chi=2\sigma_{0}^{3}k/3$ that appears in the description of the SED ${\cal S}_{\nu}$ is of the order of $10^{-3}$ for a $1$ GeV gamma-ray photon. Because the measurement of the spectrum described by eq. (5) is carried out in the cases considered here by counting the number of detected photons whose energies lie between $10^{-1}$ to $\sim 10^{3}$ GeV, the widths of the equivalent frequency bins over which each value of the flux is determined are given by $\Delta k=3\sigma_{0}^{-3}\Delta\chi/2$ with a $\Delta\chi$ whose values correspondingly range from $\sim 10^{-4}$ to $\sim 1$.
@@
While ${\hat{\kappa}}_{\rm obs}$ only contains the observed parameters of the pulsar and its emission, the value of ${\hat{\kappa}}_{\rm th}$ is determined by the physical characteristics of the magnetospheric current sheet that acts as the source of the observed emission. When, as in figure 20 of [3], the value of the colatitude $\theta_{P}$ of the observation point is sufficiently close to that of the critical angle $\theta_{P2S}$ for $|\tau_{2{\rm max}}-\tau_{2{\rm min}}|$ and hence $\sigma_{21}$ to be much smaller than unity, the right-hand side of eq. (34) is a function of the inclination angle $\alpha$ and the observer’s distance ${\hat{R}}_{P}$ only.
Equation (33) thus enables us to connect the parameters of the fitted spectra to the physical characteristics of their sources.
@@ 4.1 Characteristics of the source of the Crab pulsar’s emission
for $\Delta\chi=10^{-3}$. The dependence of ${\hat{B}}_{0}d^{2}$ on $\alpha$ implied by this relationship, in conjunction with eq. (34), is plotted in figure 5.
One can infer from the separation between the two peaks in the gamma-ray light curve of the Crab pulsar ($\sim 0.4$ cycles [16]) that the inclination angle $\alpha$ of this pulsar has a value lying in the range $60^{\circ}$ to $70^{\circ}$ (see section 5.1 of [3]). This in turn implies that the value of ${\hat{B}}_{0}d^{2}$ lies between $1.05$ and $2.61$ in this case (see figure 5), i.e. that the magnitude of the star’s magnetic field at its magnetic pole lies in the range $1.05\times 10^{12}$ to $2.61\times 10^{12}$ Gauss if we assume that the star’s radius has the value $10^{6}$ cm. These are of the same order of magnitude as the value of $B_{0}$ obtained from the conventional formula for magnetic dipole radiation (i.e. $3.79\times 10^{12}$ Gauss [21]).
The corresponding value of the privileged latitudinal direction $\theta_{P2S}$ in which the high-frequency radiation from the Crab pulsar can be observed lies between $90^{\circ}$ and $117^{\circ}$ according to figure 4. The range of values of the inclination angle $\alpha$ determines also the scale factor $j_{0}$ of the electric current density in the magnetosphere of the pulsar: from eq. (22) of [3] it follows that
@@
The above results supersede those reported in section 4 of [12].
@@ 4.2 Characteristics of the source of the Vela pulsar’s emission
Properties of the central neutron star of the Vela pulsar and its magnetosphere are more accurately reflected in the parameters of its spectrum over the photon energies $0.1$ to $10$ GeV, i.e. the interval that encompasses the majority of the data points plotted in figure 2. Insertion of the value of the fit parameter $\kappa$ given in eq. (19), the period ($0.089$ s) and the distance ($287$ pc) of this pulsar in eq. (35), together with ${\hat{\kappa}}_{\rm obs}={\hat{\kappa}}_{\rm th}$, results in
@@
for $\Delta\chi=10^{-3}$. Thus the dependence of ${\hat{B}}_{0}d^{2}$ on $\alpha$ is given by a shifted version of the curve in figure 5: one that is vertically lowered by 0.109 (cf. eqs. (36) and (38)).
@@
That the main peaks of the gamma-ray light curve of the Vela pulsar are separated by $\sim 0.43$ cycles [2] is consistent with an inclination angle close to $70^{\circ}$ (see section 5.1 of [3]).